A water is pumped at a rate of 24 cm^3/s through a 0.50 cm radius pipe on the main floor of a house to a 0.35 cm radius pipe in a solar hot water collector 4 m higher on the roof. If the pressure in the pipe on the roof is 1.2x10^5 Pa, what is the pressure in the larger pipe on the floor?


1
Expert's answer
2021-05-27T08:31:53-0400

"p_1+\rho gh+0.5rho v_1^2=p_2+0.5rho v_2^2\\"

"1.2\cdot10^5+(1000)(9.8)(4)=p_2+0.5(1000)(0.306^2-0.624^2)\\p_2=1.6\cdot10^5\ Pa"